[figure: 9.7 cm]

SUPERCONDUCTIVITY IN THE HUBBARD MODEL NEAR THE MOTT TRANSITION

Tadeusz Domański Institute of Physics, M. Curie Skłodowska University, 20-031 Lublin, Poland

Within the Hubbard model we analyse d-pairing superconductivity induced by a weak intersite attraction in a presence of a strong Coulomb onsite repulsion U, which may eventually drive a system towards the Mott transition. Effect of the local interaction is investigated using the dynamical mean field equations. Such a selfconsistent treatment is known to be exact in the limit of infinite spatial dimensions but it also represents a good approximate method for studying the finite (e.g. 2D) systems. We discuss the single particle spectrum of a normal and superconducting phases for several values of U. Other related quantities, like the optical conductivity and Pauli susceptibility show up some peculiarities which may be relevant to a description of the high T_{C} materials.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : T. Domański

Address for correspondence : Institute of Physics, M. Curie Skłodowska University, 20-031 Lublin, Poland

Email address : doman@tytan.umcs.lublin.pl